OpenStudy (anonymous):

Solve each system by graphing. x – y = 1 x + y = -7 (4, -3) (-3, -4) (-4, 3) (3, -10)

OpenStudy (anonymous):

Does anyone know the answer? :/

OpenStudy (anonymous):

(-3,-4) -3 - (-4)= 1 or -3+4=1 (because a negative and negative are a positive) and -3+(-4)= -7

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

^^ No problem

OpenStudy (anonymous):

:)